# Performance bounds for priority-based stochastic coflow scheduling

### v1

@@ 5.1 Small instances
4 4 1.37 0.14 1.25 1.46 1.66 0.13 1.57 1.75 1.92 0.24 1.77 2.09
8 4 1.57 0.15 1.47 1.69 1.72 0.16 1.63 1.82 2.30 0.31 2.12 2.52
8 8 1.48 0.11 1.41 1.56 1.68 0.11 1.61 1.76 2.65 0.29 2.45 2.87
16 8 1.60 0.08 1.55 1.66 1.71 0.10 1.65 1.78 3.07 0.27 2.91 3.26
16 16 1.57 0.07 1.53 1.62 1.70 0.09 1.64 1.76 3.74 0.26 3.56 3.91
32 16 1.60 0.06 1.55 1.65 1.66 0.07 1.61 1.71 4.01 0.31 3.81 4.17
Table 6: $C^{ALG}/C^{LP}$ for flow sizes following a Normal distribution with $CV=2$.
(a) $CV=0.5$
(b) $CV=2$
Figure 3: Histograms of the ratios $C^{ALG}/C^{LP}$ for CL, NC, RR and RO for $100$ random problem instances with $L=32$ ports and $n=16$ coflows. Flow sizes follow a Normal distribution with either (a) $CV=0.5$ or (b) $CV=2$.
@@
Figure 4: Average values of $C^{ALG}/C^{LP}$ for flow sizes following a Normal distribution with $CV=0.5$ (left), $CV=1$ (middle) or $CV=2$ (right). The number of ports $L$ and the number $n$ of coflows are indicated on the $x$-axis as $L.n$.
@@
The statistical results obtained for a Pareto distribution of flow sizes are reported in Tables 7, 8 and 9 for $CV=\frac{1}{2}$, $CV=1$ and $CV=2$, respectively. Figures 5(a) and 5(b) show the statistical distributions of the ratios $C^{ALG}/C^{LP}$ obtained for $L=32$ ports, $n=16$ coflows and $CV=\frac{1}{2}$ or $CV=2$, respectively.
As for the gamma and normal distributions, we observe that the NC is near-optimal and performs almost as well as CL when the variability of flow sizes is low, and remains quite close to it when the variability of flow sizes increases. In all considered configurations, NC outperforms RR, which in most cases is worst than an random-order policy.
CL NC RR
@@
4 4 1.02 0.05 0.97 1.06 1.06 0.05 1.03 1.10 1.59 0.11 1.54 1.69
8 4 1.07 0.05 1.03 1.11 1.10 0.08 1.05 1.13 1.72 0.19 1.60 1.85
8 8 1.04 0.04 1.01 1.07 1.09 0.05 1.05 1.12 2.11 0.17 2.02 2.21
16 8 1.09 0.04 1.06 1.11 1.11 0.05 1.08 1.15 2.23 0.18 2.11 2.34
16 16 1.08 0.03 1.06 1.10 1.12 0.05 1.08 1.15 2.79 0.21 2.65 2.94
32 16 1.10 0.03 1.08 1.13 1.12 0.04 1.08 1.15 2.90 0.24 2.75 3.02
Table 7: $C^{ALG}/C^{LP}$ for flow sizes following a Pareto distribution with $CV=0.5$.
CL NC RR
@@
4 4 0.99 0.05 0.94 1.03 1.08 0.06 1.03 1.11 1.52 0.11 1.47 1.62
8 4 1.08 0.06 1.03 1.13 1.14 0.09 1.08 1.19 1.69 0.20 1.56 1.82
8 8 1.04 0.05 1.01 1.08 1.13 0.06 1.10 1.18 2.03 0.17 1.96 2.13
16 8 1.10 0.04 1.08 1.13 1.17 0.06 1.13 1.21 2.19 0.18 2.05 2.29
16 16 1.09 0.04 1.07 1.11 1.17 0.06 1.13 1.21 2.71 0.20 2.57 2.85
32 16 1.12 0.04 1.09 1.15 1.18 0.05 1.14 1.21 2.85 0.24 2.69 2.96
Table 8: $C^{ALG}/C^{LP}$ for flow sizes following a Pareto distribution with $CV=1$.
CL NC RR
@@
4 4 0.98 0.06 0.94 1.02 1.10 0.07 1.06 1.13 1.48 0.11 1.41 1.58
8 4 1.08 0.07 1.03 1.13 1.17 0.10 1.11 1.22 1.66 0.20 1.53 1.79
8 8 1.03 0.06 1.01 1.07 1.17 0.08 1.12 1.22 1.98 0.17 1.90 2.08
16 8 1.12 0.05 1.09 1.16 1.23 0.07 1.19 1.28 2.17 0.18 2.03 2.28
16 16 1.10 0.04 1.08 1.13 1.24 0.06 1.19 1.27 2.66 0.20 2.53 2.80
32 16 1.14 0.05 1.11 1.18 1.24 0.06 1.21 1.29 2.82 0.24 2.67 2.93
Table 9: $C^{ALG}/C^{LP}$ for flow sizes following a Pareto distribution with $CV=2$.
(a) $CV=0.5$
(b) $CV=2$
Figure 5: Histograms of the ratios $C^{ALG}/C^{LP}$ for CL, NC, RR and RO for $100$ random problem instances with $L=32$ ports and $n=16$ coflows. Flow sizes follow a Pareto distribution with either (a) $CV=0.5$ or (b) $CV=2$.
@@
Figure 6: Average values of the ratios $C^{ALG}/C^{LP}$ for flow sizes following a Pareto distribution with $CV=0.5$ (left), $CV=1$ (middle) or $CV=2$ (right). The number of ports $L$ and the number $n$ of coflows are indicated on the $x$-axis as $L.n$.
@@
Those results are completely different from those obtained for synthetic distribution of flow sizes. For all considered configurations, NC performs poorly as compared to CL, and is even significantly worse than RR. The ratio $C^{NC}/C^{LP}$ is also much more variable than for synthetic distributions. For small values of $L$ and $n$, both CL and RR provide average CCT which are lower than $C^{LP}$. We remark that RR can perform better than the lower bound of LP since it falls outside the class of order-based policies. The LP bound is valid only for policies that are within this class.
@@
In conclusion, for small instances and low variability flow-sizes, the proposed NC policy is close to optimal and is better than RR. However, for high variability flow-sizes, RR outperforms NC and can even be better than the lower bound obtained from LP.
@@ 5.2 Large instances
We now consider larger problem instances for which the lower bound $C^{LP}$ cannot be computed. As the clairvoyant offline optimum is no more available, we choose as reference the clairvoyant Sincronia (CL), and present the ratios $C^{ALG}/C^{CL}$ obtained with two different algorithms: the non-clairvoyant Sincronia-based priority policy (NC) and the round-robin rate allocation (RR). As before, we average over $1000$ random traffic realizations for each problem instance. We present below the results obtained for a gamma distribution of flow sizes, as well as for flow sizes distributed as in the Facebook trace.
@@
The mean, standard deviation as well as the first and third quartiles of the ratios $C^{ALG}/C^{CL}$ are reported in Tables 11, 12 and 13 for flow sizes following a Gamma distribution with $CV=0.2$, $CV=1$ and $CV=5$, respectively. For $CV=0.2$ and $CV=1$, we observe that for all considered configurations of the numbers of ports and coflows, NC performs almost as well as CL on average, and that the variability of the ratio $C^{NC}/C^{CL}$ is quite low. We also observe that the ratio $C^{RR}/C^{CL}$ has a much higher mean and standard deviation, which shows that the round-robin allocation performs poorly as compared to our NC policy when the variability of flow sizes is low.
On the other hand, for $CV=5$ (that is, flow sizes with high variability) and a switch with fewer and equal to $40$ ports, the average performance of RR is better than but close to that of NC. However, for larger switches with $80$ ports, RR becomes worse than NC. This observation is similar to that for small instances and $CV=2$ (see Table 3) for which RR was better for smaller instances but worse for larger ones.
NC RR
$L$ $n$ mean std Q1 Q3 mean std Q1 Q3
20 10 1.00 0.02 1.00 1.01 2.29 0.32 2.13 2.49
20 20 1.01 0.02 1.00 1.01 2.82 0.26 2.64 2.98
20 30 1.01 0.02 1.00 1.02 3.13 0.27 2.95 3.30
40 20 1.00 0.02 0.99 1.01 2.84 0.41 2.55 3.15
40 40 1.00 0.02 0.99 1.01 3.52 0.30 3.34 3.72
40 60 1.01 0.01 1.00 1.02 3.97 0.27 3.81 4.15
80 40 1.01 0.02 0.99 1.02 3.78 0.33 3.58 4.04
80 80 1.01 0.02 1.00 1.02 4.62 0.32 4.41 4.87
80 120 1.00 0.01 1.00 1.01 5.11 0.33 4.91 5.33
@@
Table 11: $C^{ALG}/C^{CL}$ for flow sizes following a Gamma distribution with $CV=0.2$.
NC RR
$L$ $n$ mean std Q1 Q3 mean std Q1 Q3
20 10 1.06 0.04 1.03 1.08 1.94 0.20 1.82 2.10
20 20 1.08 0.03 1.06 1.10 2.37 0.18 2.29 2.51
20 30 1.09 0.03 1.07 1.11 2.64 0.16 2.54 2.74
40 20 1.05 0.02 1.04 1.06 2.36 0.30 2.16 2.58
40 40 1.06 0.02 1.04 1.07 2.94 0.22 2.82 3.08
40 60 1.07 0.01 1.06 1.08 3.30 0.21 3.17 3.44
80 40 1.04 0.02 1.03 1.05 3.13 0.21 3.02 3.27
80 80 1.05 0.02 1.04 1.06 3.83 0.24 3.67 4.02
80 120 1.05 0.01 1.04 1.06 4.24 0.25 4.08 4.40
@@
Table 12: $C^{ALG}/C^{CL}$ for flow sizes following a Gamma distribution with $CV=1$.
NC RR
$L$ $n$ mean std Q1 Q3 mean std Q1 Q3
20 10 1.53 0.13 1.44 1.62 1.24 0.06 1.19 1.28
20 20 1.80 0.13 1.71 1.86 1.40 0.06 1.37 1.45
20 30 1.94 0.13 1.85 2.01 1.53 0.06 1.50 1.57
40 20 1.49 0.10 1.43 1.57 1.33 0.07 1.28 1.38
40 40 1.66 0.09 1.60 1.72 1.56 0.06 1.53 1.61
40 60 1.75 0.08 1.70 1.80 1.70 0.06 1.67 1.74
80 40 1.37 0.06 1.32 1.40 1.52 0.07 1.47 1.57
80 80 1.52 0.07 1.47 1.56 1.80 0.08 1.75 1.86
80 120 1.60 0.07 1.54 1.64 1.97 0.08 1.93 2.02
@@
Table 13: $C^{ALG}/C^{CL}$ for flow sizes following a Gamma distribution with $CV=5$.
Figure 9 shows the histograms of the ratios $C^{ALG}/C^{CL}$ obtained for the $100$ random problem instances for a configuration with $80$ ports and $120$ coflows when $CV=0.2$ (left) and when $CV=5$ (right). It is clear that NC performs quite well in both cases, whereas RR performs poorly for low variability flow sizes.
(a) $CV=0.2$
(b) $CV=5$
Figure 9: Histograms of the ratios $C^{ALG}/C^{CL}$ for NC and RR for $100$ random problem instances with $L=80$ ports and $n=120$ coflows. Flow sizes follow a Gamma distribution with either (a) $CV=0.2$ or (b) $CV=5$.
@@
Figure 10: Average values of $C^{ALG}/C^{CL}$ for flow sizes following a Gamma distribution with $CV=0.2$ (left), $CV=1$ (middle) or $CV=5$ (right). The numbers of ports $L$ and of coflows $n$ are indicated on the $x$-axis as $L.n$.
@@
The statistical results for flow size distribution derived from the empirical distribution of flow sizes in the Facebook trace is reported in Table 14, and we show the specific distributions of the ratios $C^{ALG}/C^{CL}$ obtained for $L=80$ and $n=120$ in Figure 11.
For most configurations, NC performs poorly as compared to the CL, and is even significantly worse than RR. However, for instances with $80$ ports, NC outperforms (and only slightly) RR. This observation is again similar to the one for small instances (see Table 10) for NC gets closer RR in performance as the instances become larger. Again, RR is better for smaller instances because of high variability of flow sizes. An intuitive explanation for this phenomenon is as follows: as the number of flows increases, the aggregate coefficient of variation at a port will decrease thus favoring NC over RR.
NC RR
$L$ $n$ mean std Q1 Q3 mean std Q1 Q3
20 10 1.70 0.12 1.62 1.76 1.25 0.06 1.21 1.30
20 20 2.15 0.14 2.05 2.24 1.45 0.06 1.42 1.50
20 30 2.43 0.16 2.29 2.52 1.61 0.06 1.57 1.65
40 20 1.65 0.10 1.57 1.71 1.36 0.08 1.31 1.41
40 40 1.98 0.11 1.92 2.05 1.62 0.07 1.58 1.66
40 60 2.17 0.11 2.09 2.24 1.80 0.07 1.75 1.84
80 40 1.57 0.07 1.52 1.62 1.56 0.09 1.49 1.61
80 80 1.81 0.08 1.75 1.86 1.86 0.09 1.80 1.92
80 120n 1.93 0.08 1.88 1.98 2.06 0.09 2.00 2.12
Table 14: Main statistical parameters of ratios $C^{ALG}/C^{CL}$ for flow sizes distributed as in the Facebook trace.
Figure 11: Histograms of the ratios $C^{ALG}/C^{CL}$ for NC and RR for $100$ random problem instances with $L=80$ ports and $n=120$ coflows. Flow sizes are distributed as in the Facebook trace.
Figure 12 provides a synthetic view of the above results. It shows the average values (over $100$ random instances) of $C^{ALG}/C^{CL}$ obtained for the non-clairvoyant Sincronia (NC) and the round-robin allocation (RR).
Figure 12: Average values of $C^{ALG}/C^{CL}$ for NC and RR for flow sizes distributed as in the Facebook trace. The numbers of ports $L$ and of coflows $n$ are indicated on the $x$-axis as $L.n$.

### v2

@@ 5.1 Small instances
4 4 17.3 1.37 0.14 1.25 1.46 1.66 0.13 1.57 1.75 1.92 0.24 1.77 2.09
8 4 20.3 1.57 0.15 1.47 1.69 1.72 0.16 1.63 1.82 2.30 0.31 2.12 2.52
8 8 20.3 1.48 0.11 1.41 1.56 1.68 0.11 1.61 1.76 2.65 0.29 2.45 2.87
16 8 22.8 1.60 0.08 1.55 1.66 1.71 0.10 1.65 1.78 3.07 0.27 2.91 3.26
16 16 22.8 1.57 0.07 1.53 1.62 1.70 0.09 1.64 1.76 3.74 0.26 3.56 3.91
32 16 25.1 1.60 0.06 1.55 1.65 1.66 0.07 1.61 1.71 4.01 0.31 3.81 4.17
Table 6: $C^{ALG}/C^{LP}$ for flow sizes following a Normal distribution with $\eta=2$.
(a) $\eta=0.5$
(b) $\eta=2$
Figure 3: Histograms of the ratios $C^{ALG}/C^{LP}$ for CL, NC, RR and RO for $100$ random problem instances with $L=32$ ports and $n=16$ coflows. Flow sizes follow a Normal distribution with either (a) $\eta=0.5$ or (b) $\eta=2$.
@@
Figure 4: Average values of $C^{ALG}/C^{LP}$ for flow sizes following a Normal distribution with $\eta=0.5$ (left), $\eta=1$ (middle) or $\eta=2$ (right). The number of ports $L$ and the number $n$ of coflows are indicated on the $x$-axis as $L.n$.
@@
The statistical results obtained for a Pareto distribution of flow sizes are reported in Tables 7, 8 and 9 for $\eta=0.5$, $\eta=1$ and $\eta=2$, respectively. Figures 5(a) and 5(b) show the statistical distributions of the ratios $C^{ALG}/C^{LP}$ obtained for $L=32$ ports, $n=16$ coflows and $\eta=0.5$ or $\eta=2$, respectively. Column UB in these tables is the theoretical upper bound on $C^{NC}/C^{LP}$ computed using (22).
As for the gamma and normal distributions, we observe that UB is pessimistic, and that NC is near-optimal and performs almost as well as CL when the variability of flow sizes is low, and remains quite close to it when the variability of flow sizes increases. In all considered configurations, NC outperforms RR, which in most cases is worst than an random-order policy.
UB CL NC $R$
@@
4 4 8.0 1.02 0.05 0.97 1.06 1.06 0.05 1.03 1.10 1.59 0.11 1.54 1.69
8 4 9.7 1.07 0.05 1.03 1.11 1.10 0.08 1.05 1.13 1.72 0.19 1.60 1.85
8 8 9.7 1.04 0.04 1.01 1.07 1.09 0.05 1.05 1.12 2.11 0.17 2.02 2.21
16 8 12.0 1.09 0.04 1.06 1.11 1.11 0.05 1.08 1.15 2.23 0.18 2.11 2.34
16 16 12.0 1.08 0.03 1.06 1.10 1.12 0.05 1.08 1.15 2.79 0.21 2.65 2.94
32 16 15.3 1.10 0.03 1.08 1.13 1.12 0.04 1.08 1.15 2.90 0.24 2.75 3.02
Table 7: $C^{ALG}/C^{LP}$ for flow sizes following a Pareto distribution with $\eta=0.5$.
UB CL NC $R$
@@
4 4 12.0 0.99 0.05 0.94 1.03 1.08 0.06 1.03 1.11 1.52 0.11 1.47 1.62
8 4 15.3 1.08 0.06 1.03 1.13 1.14 0.09 1.08 1.19 1.69 0.20 1.56 1.82
8 8 15.3 1.04 0.05 1.01 1.08 1.13 0.06 1.10 1.18 2.03 0.17 1.96 2.13
16 8 20.0 1.10 0.04 1.08 1.13 1.17 0.06 1.13 1.21 2.19 0.18 2.05 2.29
16 16 20.0 1.09 0.04 1.07 1.11 1.17 0.06 1.13 1.21 2.71 0.20 2.57 2.85
32 16 26.6 1.12 0.04 1.09 1.15 1.18 0.05 1.14 1.21 2.85 0.24 2.69 2.96
Table 8: $C^{ALG}/C^{LP}$ for flow sizes following a Pareto distribution with $\eta=1$.
UB CL NC $R$
@@
4 4 20.0 0.98 0.06 0.94 1.02 1.10 0.07 1.06 1.13 1.48 0.11 1.41 1.58
8 4 26.6 1.08 0.07 1.03 1.13 1.17 0.10 1.11 1.22 1.66 0.20 1.53 1.79
8 8 26.6 1.03 0.06 1.01 1.07 1.17 0.08 1.12 1.22 1.98 0.17 1.90 2.08
16 8 36.0 1.12 0.05 1.09 1.16 1.23 0.07 1.19 1.28 2.17 0.18 2.03 2.28
16 16 36.0 1.10 0.04 1.08 1.13 1.24 0.06 1.19 1.27 2.66 0.20 2.53 2.80
32 16 49.2 1.14 0.05 1.11 1.18 1.24 0.06 1.21 1.29 2.82 0.24 2.67 2.93
Table 9: $C^{ALG}/C^{LP}$ for flow sizes following a Pareto distribution with $\eta=2$.
(a) $\eta=0.5$
(b) $\eta=2$
Figure 5: Histograms of the ratios $C^{ALG}/C^{LP}$ for CL, NC, RR and RO for $100$ random problem instances with $L=32$ ports and $n=16$ coflows. Flow sizes follow a Pareto distribution with either (a) $\eta=0.5$ or (b) $\eta=2$.
@@
Figure 6: Average values of the ratios $C^{ALG}/C^{LP}$ for flow sizes following a Pareto distribution with $\eta=0.5$ (left), $\eta=1$ (middle) or $\eta=2$ (right). The number of ports $L$ and the number $n$ of coflows are indicated on the $x$-axis as $L.n$.
@@
Those results are completely different from those obtained for synthetic distribution of flow sizes. For all considered configurations, NC performs poorly compared to CL, and is even significantly worse than RR. The ratio $C^{NC}/C^{LP}$ is also much more variable than for synthetic distributions. For small values of $L$ and $n$, both CL and RR provide average CCT which are lower than $C^{LP}$. We remark that RR can perform better than the lower bound of LP since it falls outside the class of order-based policies. The LP bound is valid only for policies that are within this class.
@@
In conclusion, for small instances and low variability of flow-sizes, the proposed NC policy is close to optimal and is better than RR. However, for high variability of flow-sizes, RR outperforms NC and can even be better than the lower bound obtained from LP.
@@ 5.2 Large instances
We now consider larger problem instances for which the lower bound $C^{LP}$ cannot be computed. As the clairvoyant offline optimum is no more available, we choose as reference CL, and present the ratios $C^{ALG}/C^{CL}$ obtained with three different algorithms: the non-clairvoyant Sincronia-based priority policy (NC), the round-robin rate allocation (RR ), and $\mathtt{Philae}$ which is a closely related non-clairvoyant scheduling approach proposed by A. Jajoo et al. in [JHL2022] (see also [JHLD2023]). Instead of assuming known probability distributions of flow sizes inferred from historical data, $\mathtt{Philae}$ pre-schedules sampled flows (called pilot flows) of each coflow and uses their sizes to estimate the average flow size of the coflow. Learned coflow sizes are then used to transmit flows according to a priority-based policy. This pilot-flow-based scheduling method was shown to be empirically effective in [JHL2022].
As there is no publicly available implementation of $\mathtt{Philae}$, we implemented our own version. In this simplified version, it is assumed that mean flow sizes are known by the coflow scheduler, so there is no need to sample pilot flows and no estimation error. Exact mean flow sizes are used to assign coflows to priority queues according to a contention-and-length-based policy (see policy D in [JHL2022]). Our implementation uses the same default parameters as those used for $\mathtt{Aalo}$ and $\mathtt{Philae}$ in [JHL2022] (that is, there are $K=10$ priority queues, with $Q_{0}^{hi}=10$ MB and $E=10$). As in in [JHL2022], the bandwidth of a port is shared between the priority queues according to a weighted sharing mechanism and the weights assigned to individual priority queues decrease exponentially by a factor of $10$.
As before, we average over $1000$ random traffic realizations for each problem instance. We present below the results obtained for a gamma distribution of flow sizes, as well as for flow sizes distributed as in the Facebook trace. The theoretical upper bound (UB) is not shown here since it was already quite pessimistic for small instances.
@@
We consider three different coefficients of variations: $\eta=0.2$, $\eta=1$, and $\eta=5$. The mean, standard deviation as well as the first and third quartiles of the ratios $C^{ALG}/C^{CL}$ are reported in Table 11 (resp. Table 12 and Table 13) for flow sizes following a gamma distribution with $\eta=0.2$ (resp. $\eta=1$ and $\eta=5$). For $\eta=0.2$ and $\eta=1$, we observe that for all considered configurations of the numbers of ports and coflows, NC performs almost as well as CL on average, and that the variability of the ratio $C^{NC}/C^{CL}$ is quite low. We also observe that the ratio $C^{RR}/C^{CL}$ has a much higher mean and standard deviation, which shows that RR performs poorly as compared to our NC policy when the variability of flow sizes is low. It is also interesting to note that the results obtained with $\mathtt{Philae}$ are better than those with RR but worse than that of NC. The weighted scheme of $\mathtt{Philae}$ thus outperforms the unweighted RR for small values of $\eta$.
@@
Table 11: $C^{ALG}/C^{CL}$ for flow sizes following a Gamma distribution with $\eta=0.2$ and large instances.
NC $R$ PH
$L$ $n$ mean std Q1 Q3 mean std Q1 Q3 mean std Q1 Q3
20 10 1.01 0.02 1.00 1.01 2.39 0.25 2.24 2.53 2.22 0.31 2.03 2.45
20 20 1.01 0.02 1.00 1.01 2.86 0.22 2.72 2.98 1.98 0.16 1.87 2.09
20 30 1.01 0.02 1.00 1.01 3.21 0.22 3.06 3.38 2.41 0.18 2.29 2.52
20 40 1.01 0.01 1.00 1.02 3.38 0.21 3.22 3.55 2.74 0.18 2.63 2.87
20 60 1.01 0.01 1.00 1.02 3.57 0.19 3.44 3.69 3.14 0.14 3.05 3.23
40 20 1.00 0.01 1.00 1.01 2.95 0.28 2.78 3.09 2.33 0.19 2.19 2.46
40 40 1.01 0.01 1.00 1.02 3.44 0.23 3.28 3.59 3.08 0.18 2.94 3.20
40 60 1.01 0.01 1.00 1.02 3.97 0.27 3.81 4.15 2.92 0.44 2.56 3.30
40 80 1.01 0.01 1.00 1.02 3.83 0.22 3.67 3.96 2.30 0.27 2.11 2.42
@@
Table 12: $C^{ALG}/C^{CL}$ for flow sizes following a Gamma distribution with $\eta=1$ and large instances.
NC $R$ PH
$L$ $n$ mean std Q1 Q3 mean std Q1 Q3 mean std Q1 Q3
20 10 1.06 0.04 1.03 1.08 2.12 0.16 2.04 2.22 2.00 0.22 1.87 2.15
20 20 1.07 0.03 1.04 1.08 2.56 0.14 2.47 2.65 1.88 0.14 1.78 1.98
20 30 1.07 0.02 1.06 1.09 2.86 0.15 2.77 2.97 2.28 0.16 2.20 2.41
20 40 1.08 0.02 1.06 1.09 3.02 0.14 2.93 3.13 2.59 0.15 2.51 2.69
20 60 1.08 0.02 1.06 1.09 3.21 0.12 3.13 3.30 2.94 0.10 2.88 3.01
40 20 1.04 0.03 1.02 1.05 2.72 0.20 2.60 2.83 2.24 0.18 2.11 2.35
40 40 1.05 0.02 1.03 1.06 3.18 0.18 3.05 3.32 2.93 0.15 2.82 3.05
40 60 1.05 0.02 1.04 1.06 3.40 0.18 3.29 3.52 2.79 0.40 2.46 3.14
40 80 1.05 0.02 1.04 1.06 3.58 0.17 3.45 3.70 2.20 0.23 2.05 2.28
@@
Table 13: $C^{ALG}/C^{CL}$ for flow sizes following a Gamma distribution with $\eta=5$ and large instances.
NC $R$ PH
$L$ $n$ mean std Q1 Q3 mean std Q1 Q3 mean std Q1 Q3
20 10 1.47 0.12 1.39 1.55 1.37 0.06 1.34 1.41 1.36 0.05 1.33 1.39
20 20 1.62 0.10 1.55 1.69 1.59 0.06 1.56 1.64 1.56 0.07 1.51 1.62
20 30 1.70 0.10 1.62 1.74 1.73 0.06 1.70 1.77 1.72 0.05 1.70 1.75
20 40 1.73 0.09 1.66 1.78 1.84 0.04 1.81 1.87 1.84 0.04 1.81 1.86
20 60 1.76 0.09 1.70 1.82 1.98 0.04 1.95 2.01 1.99 0.04 1.96 2.01
40 20 1.31 0.06 1.26 1.35 1.78 0.06 1.75 1.83 1.71 0.09 1.64 1.78
40 40 1.35 0.05 1.31 1.39 2.09 0.06 2.04 2.13 2.06 0.07 2.02 2.11
40 60 1.36 0.04 1.33 1.38 2.28 0.05 2.25 2.32 2.04 0.16 1.92 2.18
40 80 1.36 0.03 1.34 1.39 2.41 0.04 2.38 2.44 1.81 0.06 1.77 1.82
On the other hand, for $\eta=5$ (that is, flow sizes with high variability), a switch with $20$ ports and $10$ or $20$ coflows, the average performance of RR is better than but close to that of NC. However, for a higher number of coflows or larger switches with $40$ ports, RR becomes worse than NC. This observation is similar to that for small instances and $\eta=2$ (see Table 3) for which RR was better for smaller instances but worse for larger ones. The performance of $\mathtt{Philae}$ is close to that of RR. This is probably due to default configuration that we used for priority queues and for which many coflows fall in the same queue.
@@
(a) $C^{NC}/C^{CL}$ and $C^{RR}/C^{CL}$ with $\eta=0.2$ (left), $\eta=1$ (middle) or $\eta=5$ (right).
(b) $C^{NC}/C^{CL}$ and $C^{PH}/C^{CL}$ with $\eta=0.2$ (left), $\eta=1$ (middle) or $\eta=5$ (right).
Figure 9: Average values of the ratios $C^{ALG}/C^{CL}$ for gamma distribution of flow sizes with different coefficient of variations. Large instances. The numbers of ports $L$ and of coflows $n$ are indicated on the $x$-axis as $L.n$.
@@
The statistical results for flow size distribution derived from the empirical distribution of flow sizes in the Facebook trace is reported in Table 14, and we show the specific distributions of the ratios $C^{ALG}/C^{CL}$ obtained for $L=40$ and $n=80$ in Figure 10.
For instances with $20$ ports, NC performs poorly compared to CL, and is even significantly worse than RR. However, we observe that NC gets closer to RR for larger instances with $40$ ports. This observation is similar to the one for small instances (see Table 10): NC gets closer RR in performance as the instances become larger. We believe that RR is better for smaller instances because of high variability of flow sizes. An intuitive explanation for this phenomenon is as follows: as the number of flows increases, the aggregate coefficient of variation at a port will decrease thus favoring NC over RR. The performance of Philae is marginally worse than that of RR. Since there are several parameters to be set in Philae, it is possible that the performance will depend on the specific choice of the parameters. The results presented here are for the values mentioned previously.
One avenue for further research would be to conceive adaptive algorithms that choose between RR and NC based on the variability of the flow sizes.
Table 14: $C^{ALG}/C^{CL}$ for flow sizes distributed as in the Facebook trace and large instances.
NC PH $R$
$L$ $n$ mean std Q1 Q3 mean std Q1 Q3 mean std Q1 Q3
20 10 1.83 0.13 1.74 1.90 1.34 0.10 1.29 1.36 1.32 0.06 1.28 1.36
20 20 2.32 0.15 2.21 2.38 1.82 0.11 1.78 1.87 1.56 0.06 1.53 1.60
20 30 2.54 0.14 2.43 2.62 2.00 0.06 1.97 2.05 1.72 0.05 1.69 1.76
20 40 2.71 0.13 2.59 2.78 2.11 0.07 2.06 2.15 1.86 0.05 1.83 1.89
20 60 2.90 0.15 2.78 3.02 2.24 0.07 2.19 2.28 2.03 0.05 2.00 2.07
40 20 1.99 0.14 1.89 2.04 1.67 0.07 1.62 1.71 1.54 0.05 1.51 1.57
40 40 2.30 0.14 2.20 2.37 1.91 0.05 1.88 1.94 1.84 0.04 1.81 1.87
40 60 2.44 0.11 2.35 2.49 2.13 0.06 2.10 2.17 2.04 0.04 2.02 2.07
40 80 2.52 0.12 2.43 2.59 2.34 0.07 2.29 2.39 2.18 0.04 2.16 2.21
Figure 10: Histograms of the ratios $C^{ALG}/C^{CL}$ for $100$ random problem instances with $L=40$ ports and $n=80$ coflows. Flow sizes are distributed as in the Facebook trace.